OpenStudy (anonymous):

solving system of equations. -2x+3y=8 3x-y=-5

OpenStudy (anonymous):

you can use substitution, or elimination. i personally prefer substitution.

OpenStudy (anonymous):

do you know how to use subtitution?

OpenStudy (anonymous):

thanks for replying an yes but after i substitute 3x-y=-5,... 3x=-5+y ...x=-5/3+y/3 after this im stuck like a red stop sign

OpenStudy (anonymous):

just waiting for a green light lol @WasTim

OpenStudy (anonymous):

ah okay great start, i see you modified the second equation. now notice something. x is now in terms of y. so if we sub this back into the first equation, we should be able to solve for y.

OpenStudy (anonymous):

so sub x=-5/3+y/3 into the first equation, as x.

OpenStudy (anonymous):

yes but that where im stuck like i just dont know what to do but i know what to do at the same time its confusing because its fraction type

OpenStudy (anonymous):

okay, notice we found x right?: x=-5/3+y/3 cool. notice the first equation we havent messed with: -2x+3y=8 i notice it has an x.. well we found an x before.. so lets sub that in -2(-5/3+y/3)+3y=8 can you simplify and solve for y now?

OpenStudy (anonymous):

yes thank u soo much its much clearer now but im gonna do it as soon as i play a game of call of duty ...... thanks again

OpenStudy (anonymous):

the problems half done lol..

OpenStudy (anonymous):

yea i the problem is half done kinda figured that out right now so i got after solving for y i got -2(-5/3+y/3)+3y=8.........3.3-0.66y+3y=8......2.34y=8-3.3... 2.34y=4.7 y=4.7/2.34 i got y=2.008 so so far is it correct @WasTim